# Drop rebound at low Weber number

### v1

@@ Abstract
We study the rebound of drops impacting non-wetting substrates at low Weber number $W\kern-1.49994pte\,$ through experiment, direct numerical simulation, and reduced-order modeling. Submillimeter-sized drops are normally impacted onto glass slides coated with a thin viscous film that allows them to rebound without contact line formation. Experiments are performed with various drop viscosities, sizes, and impact velocities, and we directly measure metrics pertinent to spreading, retraction, and rebound using high-speed imaging. We complement experiments with direct numerical simulation and a fully predictive reduced-order model that applies natural geometric and kinematic constraints to simulate the drop shape and dynamics using a spectral method. At low $W\kern-1.49994pte\,$, drop rebound is characterized by a weaker dependence of the coefficient of restitution on $W\kern-1.49994pte\,$ than in the more commonly studied high-$W\kern-1.49994pte\,$ regime, with nearly $W\kern-1.49994pte\,$-independent rebound in the inertio-capillary limit, and an increasing contact time as $W\kern-1.49994pte\,$ decreases. Drops with higher viscosity or size interact with the substrate longer, have a lower coefficient of restitution, and stop bouncing sooner, in good quantitative agreement with our reduced-order model. In the inertio-capillary limit, low $W\kern-1.49994pte\,$ rebound is characterized by nearly symmetric spreading and retraction phases, which corresponds to a coefficient of restitution that is near unity. Increasing $W\kern-1.49994pte\,$ or viscosity breaks this symmetry, coinciding with a drop in the coefficient of restitution and an increased dependence on $W\kern-1.49994pte\,$. Lastly, the maximum drop deformation and spreading are related through energy arguments, providing a comprehensive framework for drop impact and rebound at low $W\kern-1.49994pte\,$.
@@ 1 Introduction
Over a century ago, Lord Rayleigh observed drops rebounding upon impact (Rayleigh, 1879) and attributed this phenomenon to an intervening air layer separating the drops and preventing coalescence (Rayleigh, 1899). Since then, experiments conducted under modified atmospheric conditions and corresponding theoretical studies have confirmed his hypothesis (Willis & Orme, 2000; Bach et al., 2004; Chubynsky et al., 2020). Meanwhile, advancements in experimental and computational techniques have enabled measurement of the minimum thickness of the air layer—on the order of 100 nm—below which van der Waals forces bridge the liquid-air interfaces and capillary forces drive coalescence (Couder et al., 2005a; de Ruiter et al., 2012; Kolinski et al., 2012; Kavehpour, 2015; Sprittles, 2024; Lewin-Jones et al., 2024). Thus, drops only rebound without contact from smooth surfaces. Although there are limited solid surfaces smooth enough to promote contactless rebound, liquid-gas interfaces are atomically smooth and ideal for rebound. For example, recent results for the rebound of a drop on a bath of the same fluid showed that the deformation of the drop and bath are dynamically coupled (Alventosa et al., 2023; Phillips & Milewski, 2024), and that vibrating the bath can lead to long-lasting bouncing and wave-propelled drop motion (Couder et al., 2005b; Moláček & Bush, 2013). If the bath height is reduced to a thin film, rebound can become substrate-independent (Hao et al., 2015; Sanjay et al., 2023b). Although drops can rebound from heated substrates (Biance et al., 2006), dry ice (Li et al., 2024), and when the drop is coated with hydrophobic powder (Aussillous & Quéré, 2006), these methods inject energy into the system or modify the drop composition, potentially altering the rebound dynamics and obfuscating the fundamental physics.
@@
The Weber number $W\kern-1.49994pte\,$ compares inertial and surface tension forces, Bond number $B\kern-0.50003pto\,$ compares gravitational and surface tension forces, and Ohnesorge number $O\kern-0.50003pth\,$ is the ratio of the inertio-capillary to visco-capillary timescale. Here, $R$ is the undeformed drop radius, $V$ is the impact velocity, $\rho$ is the liquid density, $\mu$ is the dynamic viscosity, $\sigma$ is the surface tension coefficient, and $g$ is the gravitational acceleration. Although early studies emphasized $W\kern-1.49994pte\,$ as the primary parameter controlling rebound (Foote, 1975; Anders et al., 1993; Richard & Quéré, 2000; Richard et al., 2002; Okumura et al., 2003), recent work has shown that increasing $B\kern-0.50003pto\,$ and $O\kern-0.50003pth\,$ can suppress rebound for $W\kern-1.49994pte\,>1$ (Jha et al., 2020; Sanjay et al., 2023a). For $W\kern-1.49994pte\,<1$, the role of $B\kern-0.50003pto\,$ and $O\kern-0.50003pth\,$ is less explored, and attempts to isolate the role of $W\kern-1.49994pte\,$ on rebound have been confounded by the increasing influence of contact line friction as $W\kern-1.49994pte\,$ decreases (Wang et al., 2022; Thenarianto et al., 2023). Impacting a drop on a substrate coated with a viscous film avoids contact line formation, as discussed by Gilet et al. (2008) and Gilet & Bush (2012). This technique produces substrate-independent rebound when the film thickness satisfies $h_{f}\lesssim\left(\mu_{f}/W\kern-1.49994pte\,^{2}\right)^{1/3}$, where $h_{f}$ and $\mu_{f}$ are the thickness and viscosity of the film, respectively (Hao et al., 2015). Additionally, Sanjay et al. (2023b) recently identified the competing influences of the drop Ohnesorge number $O\kern-0.50003pth\,$ and film Ohnesorge number $O\kern-0.50003pth\,_{f}=\mu_{f}/\left(\rho\sigma R\right)^{1/2}$, and found that substrate-independent rebound occurs for $O\kern-0.50003pth\,_{f}\gtrsim 0.1$ when $\Gamma<\textrm{B}\,O\kern-0.50003pth\,_{f}^{1/3}$, where $\Gamma=h_{f}/R$ and $\textrm{B}=\textrm{B}\left(O\kern-0.50003pth\,\right)$. Since $\textrm{B}\approx 0.1$ when $O\kern-0.50003pth\,\approx 0.1$, and B increases with decreasing $O\kern-0.50003pth\,$, a conservative condition for substrate-independent bouncing is $\Gamma/O\kern-0.50003pth\,_{f}^{1/3}<0.1$. Notably, this result is derived for $W\kern-1.49994pte\,>1$ and becomes less restrictive for $W\kern-1.49994pte\,<1$. By following these guidelines, a rigid substrate can be prepared that allows a drop to rebound without contact line formation or dynamic influence from the viscous film. An example of such a rebound is imaged in figure 1.
@@
Figure 1: A drop of silicone oil ($R=0.31$ mm) rebounding from a glass slide coated with a thin, viscous film ($\Gamma/O\kern-0.45003pth\,_{f}^{1/3}\approx 10^{-2}$). Here, $W\kern-1.34995pte\,=1.03$, $O\kern-0.45003pth\,=0.02$, and $B\kern-0.45003pto\,=0.04$. Subsequent images are separated by 1.44 ms and shifted 0.4 mm to the right.
@@
When a drop impacts a rigid substrate, it first spreads until reaching a maximum equatorial radius $R_{m}$ and maximum contact radius $R_{c}$. When $W\kern-1.49994pte\,\gg 1$, the dimensionless maximum equatorial radius $\beta=R_{m}/R$ and maximum contact radius $\beta_{c}=R_{c}/R$ are approximately equal, and two competing scaling laws have been proposed to describe their evolution with $W\kern-1.49994pte\,$. Clanet et al. (2004) proposed $\beta\sim W\kern-1.49994pte\,^{1/4}$ based on momentum conservation and assuming the drop deforms into a cylinder with negligible height during impact. Alternatively, the conservation of energy between the kinetic energy of the drop at impact and the surface energy at the maximum deformation of the drop predicts $\beta\sim W\kern-1.49994pte\,^{1/2}$ (Chandra & Avedisian, 1991; Bennett & Poulikakos, 1993; Okumura et al., 2003). Although additional arguments and empirical evidence favor the scaling $W\kern-1.49994pte\,^{1/2}$ (Eggers et al., 2010; Villermaux & Bossa, 2011; Laan et al., 2014; Lee et al., 2016; Josserand & Thoroddsen, 2016), extensive experimental results support the scaling $W\kern-1.49994pte\,^{1/4}$ over a wide range of $W\kern-1.49994pte\,$ (Bartolo et al., 2005; Biance et al., 2006; Tsai et al., 2011; García-Geijo et al., 2020; Zhan et al., 2021; Li et al., 2024). When viscous forces are important, these scalings break down (Rein, 1993), and describing the crossover from the capillary to viscous regime requires consideration of finite $O\kern-0.50003pth\,$ effects (or $Re=\sqrt{W\kern-1.49994pte\,}/O\kern-0.50003pth\,=\rho VR/\mu$) (Roisman, 2009; Eggers et al., 2010; Laan et al., 2014; Lee et al., 2016; Sanjay & Lohse, 2025). For $W\kern-1.49994pte\,<1$, the deformed drop resembles an oblate ellipsoid rather than a wide cylinder and $\beta\not\approx\beta_{c}$. In this case, Richard & Quéré (2000) predicted $\beta-1=\left(5/48\right)^{1/2}\,W\kern-1.49994pte\,^{1/2}$ by assuming the initial kinetic energy of the drop was converted to surface energy but did not test their prediction across a range of $W\kern-1.49994pte\,$. Recently, efforts to describe and relate $\beta$ and $\beta_{c}$ during low $W\kern-1.49994pte\,$ rebound have used drops with non-negligible $B\kern-0.50003pto\,$ and relied on fits to empirical relations (Liu et al., 2023; Li et al., 2024). Thus, a comprehensive and validated understanding of spreading for low $W\kern-1.49994pte\,$ is still lacking.
Following spreading, surface tension drives the drop to retract and can lead to rebound. When $O\kern-0.50003pth\,$ is negligible, the maximum retraction rate is independent of impact velocity, and the retraction time $t_{r}\sim t_{\sigma}$, where $t_{\sigma}=\left(\rho R^{3}/\sigma\right)^{1/2}$ is the intertio-capillary time (Bartolo et al., 2005). This result is valid for $W\kern-1.49994pte\,>1$, and since inertia also governs the spreading time $t_{s}$, the interaction time between the substrate and drop $t_{c}=t_{s}+t_{r}$, referred to as the contact time henceforth, scales as $t_{c}\sim A\,t_{\sigma}$ (Richard et al., 2002), where $A=f\left(O\kern-0.50003pth\,\right)$ (Jha et al., 2020). Alternatively, when $W\kern-1.49994pte\,<1$, the non-dimensional contact time $t_{c}/t_{\sigma}$ increases as $W\kern-1.49994pte\,$ decreases (Foote, 1975; Okumura et al., 2003). This transition in $W\kern-1.49994pte\,$ dependence signals a shift in the underlying physics. This shift also alters the $W\kern-1.49994pte\,$ dependence of the coefficient of restitution $\varepsilon$, a measure of the elasticity of the rebound. For $W\kern-1.49994pte\,>1$, Anders et al. (1993) noted that $\varepsilon$ is $W\kern-1.49994pte\,$ dependent and Biance et al. (2006) showed that $\varepsilon\sim We^{-1/2}$, which has been verified on various rigid substrates (Wang et al., 2022; Thenarianto et al., 2023). Additionally, Jha et al. (2020) showed that increasing $O\kern-0.50003pth\,$ decreases the coefficient of restitution $\varepsilon$. For $W\kern-1.49994pte\,<1$, Gilet & Bush (2012) suggested $W\kern-1.49994pte\,$-independent rebound such that $\varepsilon=f\left(O\kern-0.50003pth\,\right)$. Recent efforts to study $\varepsilon$ for low $W\kern-1.49994pte\,$ rebounds observed a peak in $\varepsilon$ as $W\kern-1.49994pte\,$ decreases, followed by a decrease to $\varepsilon=0$ (no rebound) at finite $W\kern-1.49994pte\,$ (Wang et al., 2022; Thenarianto et al., 2023). The suggested functional form of this decrease $\varepsilon\sim W\kern-1.49994pte\,^{1/2}$ is attributed to contact line friction, highlighting the need for contact line-free conditions to determine the $\varepsilon$$W\kern-1.49994pte\,$ relationship at low $W\kern-1.49994pte\,$.
@@
Herein, we present experiments and direct numerical simulation results of drop rebound at low $W\kern-1.49994pte\,$, and develop a robust reduced-order model using the KM method. The experimental apparatus, testing protocol, and image processing technique are detailed in §$\S$ 2. In §$\S$ 3 we present the formulation of our reduced-order model and describe our direct numerical simulation (DNS) framework. In §$\S$ 4, we compare our experiments and DNS results with our model predictions for the pertinent metrics describing drop rebound (§$\S$ 4.1), spreading and retraction dynamics (§$\S$ 4.2), and maximum drop deformation and spreading (§$\S$ 4.3). In §$\S$ 5 we synthesize our findings into new knowledge about drop rebound and contextualize it within the existing literature. Finally, we summarize our key findings in §$\S$ 6 and identify a number of open questions inspired by this work.
@@ 2.1 Experimental setup
An atomically smooth, non-wetting substrate was prepared by coating a standard glass slide with a thin film of viscous silicone oil (Gilet & Bush, 2012; Hao et al., 2015; Sanjay et al., 2023b). To achieve this, we first coated the slide with a thick layer of viscous silicone oil $\mu=975$ g cm-1s-1. Then, the slide was passed beneath a fixed knife edge to thin and level the film. The mean film thickness $h_{f}$ was measured by weighing the slide, and was $h_{f}=231\pm 55\ \mu$m across all experiments. This thickness corresponds to a dimensionless film thickness $\Gamma=h_{f}/R\approx 10^{0}$ and film mobility $\Gamma/O\kern-0.50003pth\,_{f}^{1/3}\approx 10^{-2}$ that is well below the $\Gamma/O\kern-0.50003pth\,_{f}^{1/3}\leq 10^{-1}$ criterion discussed earlier and thus confidently within the regime of substrate-independent drop bouncing (Sanjay et al., 2023b). After the slide was prepared it was placed in a closed environment for at least 12 hours to allow non-uniformities in the film to relax under gravity. During testing, the substrate was mounted perpendicular to the direction of gravitational acceleration to ensure the air layer between the impacting drop and substrate did not rupture prematurely due to asymmetry (Lo et al., 2017).
Figure 2: ($a$) A rendering of the experimental setup. The inset depicts the drop generation process with the evolution of the liquid-air interface (left half) illustrated with colored lines increasing with time from yellow (light) to blue (dark). ($b$) Schematic of an impacting drop and measured variables. ($c$) Height of the center of mass $h$ against time $t$ for a rebounding drop ($R=0.2$ mm), where $W\kern-1.34995pte\,=0.25$, $O\kern-0.45003pth\,=0.03$, and $B\kern-0.45003pto\,=0.02$. Dashed and dot-dashed lines are parabolic fits to the drop position during free-fall and rebound, respectively. Gray dashed lines indicate the initial contact time $t=0$ and detachment time $t=t_{c}$.
@@
Drop radius $R$ cm $0.0094-0.0967$
Impact velocity $V$ cm s-1 $1.00-74.39$
Density $\rho$ g cm-3 $0.850-0.975$
Surface tension $\sigma$ g s-2 $18.7-20.8$
Dynamic viscosity $\mu$ g cm-1 s-1 $0.0187-0.4785$
@@
Weber number $W\kern-1.49994pte\,$- $0.0008-10.6185$
Ohnesorge number $O\kern-0.50003pth\,$- $0.0139-0.7865$
Bond number $B\kern-0.50003pto\,$- $0.0037-0.4197$

### v2

@@ Abstract
We study the rebound of drops impacting non-wetting substrates at low Weber number $W\kern-1.49994pte\,$ through experiment, direct numerical simulation, and reduced-order modeling. Submillimeter-sized drops are normally impacted onto glass slides coated with a thin viscous film that allows them to rebound without contact line formation. Experiments are performed with various drop viscosities, sizes, and impact velocities, and we directly measure metrics pertinent to spreading, retraction, and rebound using high-speed imaging. We complement experiments with direct numerical simulation and a fully predictive reduced-order model that applies natural geometric and kinematic constraints to simulate the drop shape and dynamics using a spectral method. At low $W\kern-1.49994pte\,$, drop rebound is characterized by a weaker dependence of the coefficient of restitution on $W\kern-1.49994pte\,$ than in the more commonly studied high-$W\kern-1.49994pte\,$ regime, with nearly $W\kern-1.49994pte\,$-independent rebound in the inertio-capillary limit, and an increasing contact time as $W\kern-1.49994pte\,$ decreases. Drops with higher viscosity or size interact with the substrate longer, have a lower coefficient of restitution, and stop bouncing sooner, in good quantitative agreement with our reduced-order model. In the inertio-capillary limit, low $W\kern-1.49994pte\,$ rebound has nearly symmetric spreading and retraction phases and a coefficient of restitution near unity. Increasing $W\kern-1.49994pte\,$ or viscosity breaks this symmetry, coinciding with a drop in the coefficient of restitution and an increased dependence on $W\kern-1.49994pte\,$. Lastly, the maximum drop deformation and spreading are related through energy arguments, providing a comprehensive framework for drop impact and rebound at low $W\kern-1.49994pte\,$.
@@ 1 Introduction
Over a century ago, Lord Rayleigh observed drops rebounding upon impact (Rayleigh, 1879) and attributed this phenomenon to an intervening air layer separating the drops and preventing coalescence (Rayleigh, 1899). Since then, experiments conducted under modified atmospheric conditions and corresponding theoretical studies have confirmed his hypothesis (Willis & Orme, 2000; Bach et al., 2004; Chubynsky et al., 2020). Meanwhile, advancements in experimental and computational techniques have enabled measurement of the minimum thickness of the air layer—on the order of 100 nm—below which van der Waals forces bridge the liquid-air interfaces and capillary forces drive coalescence (Couder et al., 2005a; de Ruiter et al., 2012; Kolinski et al., 2012; Kavehpour, 2015; Sprittles, 2024; Lewin-Jones et al., 2024). Thus, drops only rebound without contact from smooth surfaces. Although there are limited solid surfaces smooth enough to promote contactless rebound, liquid-gas interfaces are atomically smooth and ideal for rebound. For example, recent results for the rebound of a drop on a bath of the same fluid showed that the deformation of the drop and bath are dynamically coupled (Alventosa et al., 2023; Phillips & Milewski, 2024), and that vibrating the bath can lead to long-lasting bouncing and wave-propelled drop motion (Couder et al., 2005b; Moláček & Bush, 2013). This dynamic coupling is further evidenced by the complex bubble patterns observed when an impacting drop merges with a bath (Saylor & Bounds, 2012). If the bath height is reduced to a thin film, rebound can become substrate-independent (Hao et al., 2015; Sanjay et al., 2023b). Although drops can rebound from heated substrates (Biance et al., 2006), dry ice (Li et al., 2024), and when the drop is coated with hydrophobic powder (Aussillous & Quéré, 2006), these methods inject energy into the system or modify the drop composition, potentially altering the rebound dynamics and obfuscating the fundamental physics.
@@
The Weber number $W\kern-1.49994pte\,$ compares inertial and surface tension forces, Bond number $B\kern-0.50003pto\,$ compares gravitational and surface tension forces, and Ohnesorge number $O\kern-0.50003pth\,$ is the ratio of the inertio-capillary to the viscous timescale. Here, $R$ is the undeformed drop radius, $V$ is the impact velocity, $\rho$ is the liquid density, $\mu$ is the dynamic viscosity, $\sigma$ is the surface tension coefficient, and $g$ is the gravitational acceleration. Although early studies emphasized $W\kern-1.49994pte\,$ as the primary parameter controlling rebound (Foote, 1975; Anders et al., 1993; Richard & Quéré, 2000; Richard et al., 2002; Okumura et al., 2003), recent work has shown that increasing $B\kern-0.50003pto\,$ and $O\kern-0.50003pth\,$ can suppress rebound for $W\kern-1.49994pte\,>1$ (Jha et al., 2020; Sanjay et al., 2023a). For $W\kern-1.49994pte\,<1$, the role of $B\kern-0.50003pto\,$ and $O\kern-0.50003pth\,$ is less explored, and attempts to isolate the role of $W\kern-1.49994pte\,$ on rebound have been confounded by the increasing influence of contact line friction as $W\kern-1.49994pte\,$ decreases (Wang et al., 2022; Thenarianto et al., 2023). Impacting a drop on a substrate coated with a viscous film avoids contact line formation, as discussed by Gilet et al. (2008) and Gilet & Bush (2012). This technique produces substrate-independent rebound when the film is sufficiently thin and viscous (Hao et al., 2015). To determine the criteria where drop rebound is substrate-independent, Sanjay et al. (2023b) analyzed the competing influences of the drop Ohnesorge number $O\kern-0.50003pth\,$ and film Ohnesorge number $O\kern-0.50003pth\,_{f}=\mu_{f}/\left(\rho\sigma R\right)^{1/2}$, and found that substrate-independent rebound occurs for $O\kern-0.50003pth\,_{f}\gtrsim 0.1$ when $\Gamma<\textrm{B}\,O\kern-0.50003pth\,_{f}^{1/3}$, where $\Gamma=h_{f}/R$ and $\textrm{B}=\textrm{B}\left(O\kern-0.50003pth\,\right)$. Since $\textrm{B}\approx 0.1$ when $O\kern-0.50003pth\,\approx 0.1$, and B increases with decreasing $O\kern-0.50003pth\,$, a conservative condition for substrate-independent bouncing is $\Gamma/O\kern-0.50003pth\,_{f}^{1/3}<0.1$. Notably, this result is derived for $W\kern-1.49994pte\,>1$ and becomes less restrictive for $W\kern-1.49994pte\,<1$. By following these guidelines, a rigid substrate can be prepared that allows a drop to rebound without contact line formation or dynamic influence from the viscous film. An example of such a rebound is imaged in figure 1.
@@
Figure 1: A drop of silicone oil ($R=0.31$ mm) rebounding from a glass slide coated with a thin, viscous film ($\Gamma/O\kern-0.46252pth\,_{f}^{1/3}\approx 10^{-2}$). Here, $W\kern-1.38742pte\,=1.03$, $O\kern-0.46252pth\,=0.02$, and $B\kern-0.46252pto\,=0.04$. Subsequent images are separated by 1.44 ms and shifted 0.4 mm to the right.
@@
When a drop impacts a rigid substrate, it first spreads until reaching a maximum equatorial radius $R_{m}$ and maximum contact radius $R_{c}$. When $W\kern-1.49994pte\,\gg 1$, the dimensionless maximum equatorial radius $\beta=R_{m}/R$ and maximum contact radius $\beta_{c}=R_{c}/R$ are approximately equal, and two competing scaling laws have been proposed to describe their evolution with $W\kern-1.49994pte\,$. Clanet et al. (2004) proposed $\beta\sim W\kern-1.49994pte\,^{1/4}$ based on momentum conservation and assuming the drop deforms into a cylinder with negligible height during impact. Alternatively, the conservation of energy between the kinetic energy of the drop at impact and the surface energy at the maximum deformation of the drop predicts $\beta\sim W\kern-1.49994pte\,^{1/2}$ (Chandra & Avedisian, 1991; Bennett & Poulikakos, 1993; Okumura et al., 2003). Although additional arguments and empirical evidence favor the scaling $W\kern-1.49994pte\,^{1/2}$ (Eggers et al., 2010; Villermaux & Bossa, 2011; Laan et al., 2014; Lee et al., 2016; Josserand & Thoroddsen, 2016), extensive experimental results support the scaling $W\kern-1.49994pte\,^{1/4}$ over a wide range of $W\kern-1.49994pte\,$ (Bartolo et al., 2005; Biance et al., 2006; Tsai et al., 2011; Garc´ıa-Geijo et al., 2020; Zhan et al., 2021; Li et al., 2024). When viscous forces are important, these scalings break down (Rein, 1993), and describing the crossover from the capillary to viscous regime requires consideration of finite $O\kern-0.50003pth\,$ effects (or $Re=\sqrt{W\kern-1.49994pte\,}/O\kern-0.50003pth\,=\rho VR/\mu$) (Roisman, 2009; Eggers et al., 2010; Laan et al., 2014; Lee et al., 2016; Sanjay & Lohse, 2025). For $W\kern-1.49994pte\,<1$, the deformed drop resembles an oblate ellipsoid rather than a wide cylinder and $\beta\not\approx\beta_{c}$. In this case, Richard & Quéré (2000) predicted $\beta-1=\left(5/48\right)^{1/2}\,W\kern-1.49994pte\,^{1/2}$ by assuming the initial kinetic energy of the drop was converted to surface energy but did not test their prediction across a range of $W\kern-1.49994pte\,$. Recently, efforts to describe and relate $\beta$ and $\beta_{c}$ during low $W\kern-1.49994pte\,$ rebound have used drops with non-negligible $B\kern-0.50003pto\,$ and relied on fits to empirical relations (Liu et al., 2023; Li et al., 2024). Thus, a comprehensive and validated understanding of spreading for low $W\kern-1.49994pte\,$ is still lacking.
[…]
@@
Herein, we present experiments and direct numerical simulation results of drop rebound at low $W\kern-1.49994pte\,$, and develop a robust reduced-order model using the KM method. The experimental apparatus, testing protocol, and image processing technique are detailed in $\S$ 2. In $\S$ 3 we present the formulation of our reduced-order model and describe our direct numerical simulation (DNS) framework. In $\S$ 4, we compare our experiments and DNS results with our model predictions for the pertinent metrics describing drop rebound ($\S$ 4.1), spreading and retraction dynamics ($\S$ 4.2), and maximum drop deformation and spreading ($\S$ 4.3). In $\S$ 5 we synthesize our findings into new knowledge about drop rebound and contextualize it within the existing literature. Finally, we summarize our key findings in $\S$ 6 and identify a number of open questions inspired by this work.
@@ 2.1 Experimental setup
[…]
Figure 2: ($a$) A rendering of the experimental setup. The inset depicts the drop generation process with the evolution of the liquid-air interface (left half) illustrated with colored lines increasing with time from yellow (light) to blue (dark). ($b$) Schematic of an impacting drop and measured variables. ($c$) Height of the center of mass $h$ against time $t$ for a rebounding drop ($R=0.2$ mm), where $W\kern-1.38742pte\,=0.25$, $O\kern-0.46252pth\,=0.03$, and $B\kern-0.46252pto\,=0.02$. Dashed and dot-dashed lines are parabolic fits to the drop position during free-fall and rebound, respectively. Gray dashed lines indicate the initial contact time $t=0$ and detachment time $t=t_{c}$.
@@
Drop radius $R$ cm 0.00940.09670.0094-0.0967
Impact velocity $V$ cm s-1 1.0074.391.00-74.39
Density $\rho$ g cm-3 0.8500.9750.850-0.975
Surface tension $\sigma$ g s-2 18.720.818.7-20.8
Dynamic viscosity $\mu$ g cm-1 s-1 0.01870.47850.0187-0.4785
@@
Weber number $W\kern-1.49994pte\,$ - 0.000810.61850.0008-10.6185
Ohnesorge number $O\kern-0.50003pth\,$ - 0.01390.78650.0139-0.7865
Bond number $B\kern-0.50003pto\,$ - 0.00370.41970.0037-0.4197